# Machine Learning needs its own Randomness Standard: Randomised Smoothing and PRNG-based attacks

### v1

@@ 3.4 Attacks, Dual_EC_DRBG and Bullrun
Traditional attacks on pseudo-random number generators have focused on determining the next number to be generated in a sequence; or more generally, given some of the numbers in a sequence, deducing some others. If the generator is used to produce cryptographic keys and also other random values that may become visible to the attacker, such a deduction may compromise an encrypted message. Most attacks aim to determine the state of the PRNG being used by the victim. Since most PRNGs are deterministic, a known state can be used to determine future numbers in the sequence [21]. Another class of attacks involves introducing a backdoor in the PRNG to make future outputs more predictable. The NSA pushed NIST to include EC-DRBG in the ANSI X9.82 standard for random number generators when the standard was first developed at the start of the 21st century. The PRNG featured in the first draft of the standard, which was published in 2004. The research community expressed concerns about EC-DRBG not being cryptographically sound by 2005 [24] – before the first official version of the standard was published [4] – leading to conjecture of an NSA backdoor [60]. Despite these concerns, EC-DBRG was adopted as the default random number generator by RSA in their BSAFE cryptography library [15]. The Snowden leaks in 2013 confirmed the existence of project Bullrun, which aimed “to covertly introduce weaknesses into the encryption standards followed by hardware and software developers worldwide" and successfully injected a backdoor into the default PRNG used for all cryptographic encryption between 2006 and 2013. This is a useful reminder of the tactics, tools, and procedures that may be used by a capable motivated adversary to carry out a backdoor attack against pseudo-random number generators.
@@ 4.1 Threat Model
It is assumed that the victim is attempting to use ML to get predictions in a safety-sensitive or security-sensitive setting. Therefore it is important to accurately gauge the robustness of every prediction, and we assume that Randomised Smoothing is being employed for this purpose. By definition, the probability that an adversarial example can be found within the certified radius obtained from randomised smoothing is low: less than 0.1% if $\alpha$ is set to $0.001$ as suggested by Cohen et al. [13] when they proposed this technique. The knowledge that the victim is employing randomised smoothing can in itself be very useful to an adversary. This was demonstrated by Cullen et al. [16], where they used this information to only search for adversarial examples with l2 distance greater than the certified radius, achieving better than the state-of-the-art success rate at finding adversarial examples against ML models using randomised smoothing. Furthermore, in their evaluation of randomised smoothing, Cohen et al. [13] found that the probability of finding adversarial examples increases rapidly as the upper bound on the l2 norm of the adversarial example set by the adversary increases beyond the certified radius, $R$. For the ImageNet dataset [17], they found that the probability of finding an adversarial example against a smoothing classifier is 0.17 at an upper bound of $1.5R$ and 0.53 at $2R$. Hence, confidence serves a good proxy for reliability of prediction, and any compromise of Randomised Smoothing will decrease its trustworthiness.
@@
The objectives of an attacker attempting the class of attacks presented in this paper are twofold, depending on whether the certified radius obtained from randomised smoothing is being manipulated to be higher or lower. A higher certified radius can make the victim believe that the prediction for a given input is more robust than it actually is, so that the victim ignores adversarial examples that exist within the spoofed radius. On the other hand, certification is costly, requiring approximately $10^{5}$ inferences from the smoothed model to certify a radius of $4\sigma$. So reducing the certified radius can cause the victim to waste time and compute resources, forcing them to generate more samples of predictions under noise to obtain the desired radius – and thus providing a service-denial attack.
@@
The setup used by the victim to perform training and certification can itself be quite complicated, leaving multiple avenues for an attacker to gain access and manipulate the noise being used to perform the sampling. Modern practices such as ML-as-a-service, and outsourcing of training and data generation to third-parties, opened a Pandora’s box of new attacks. The attacks discussed here focus on modifying the noise distribution being used, first by using a different noise function directly and later by modifying the bitstream generated by the underlying pseudo-random number generator (PRNG). Of course, where the victim has little or no control over the software and hardware used for training and certification, there are even more attack vectors to worry about. The objective of the attacks presented in the following sections is to spoof the certified radius, while escaping detection by the victim, whether by analysing the performance of the model, or by using statistical tests (discussed in more detail in Section 4.4). This class of attacks can be carried out by targeting the hardware and software layers as well. In addition, with recent developments in denoising diffusion probabilistic models [27], such attacks can be carried out in an ML-as-a-Service scenario as well. If the victim is submitting inputs to a cloud API for prediction and certification, an attacker can effectively remove the noise introduced by the victim, and replace it.
@@ 4.2 Naive noise distribution replacement
In order to check the feasibility of the attacks described in the previous section, an initial test was done by explicitly changing the distribution that is sampled during certification. The Github repository released by Cohen et al. [^6] https://github.com/Xzh0u/randomized-smoothing was used as a starting point. Cohen et al. [13] reported that the variance of distribution that additive noise is sampled from controls the trade off between the accuracy and robustness. A lower value of $\sigma$ can be used to certify smaller radii but with a high degree of accuracy, whereas a higher value of $\sigma$ is needed to certify large radii, resulting in lower accuracy. $\sigma=0.25$ was found to achieve an acceptable trade off between certification radius and accuracy on the CIFAR 10 [35] dataset, which was used to run all the experiments in this paper. Using a base classifier trained with additive noise sampled from $\mathcal{N}(0,0.25^{2})$, the naive attack swaps the noise distribution for one of the following when a certification is performed: Laplace distribution, $\mathcal{L}(0,0.25)$; Absolute value of normal distribution, $|\mathcal{N}(0,0.25^{2})|$; Uniform distribution, $\mathcal{U}(-0.25,0.25)$; and Bernoulli distribution, $\mathcal{B}(0.5)$.
This naive attack was used to demonstrate the feasibility of this class of attacks on PRNGs. The next section will describe a more sophisticated attack.
@@ 4.3 Bit-flipping PRNG attacker
The objective of the bit-flipping PRNG attacker is to modify the stream of bits being generated by the random number generator to alter the distribution when normal variates are sampled. An overview of the algorithm used to transform random integers to floats in the standard normal distribution is presented in Section 3.3. Taking a 64-bit random integer, the rightmost 8 bits determine the rectangle in which the normally distributed random number will fall. These are called the index bits. The 9th bit is the sign bit. Finally, the remaining 55 bits are transformed into a floating point number falling within the limits of the uniformly distributed rectangle chosen by the index bits. These are called the distribution bits. The following attacks modify one of these three categories of bits to alter the resulting distribution. Altering a distribution can be done by either introducing kurtosis or skewness. Kurtosis is a measure of the tailedness of a distribution whereas skewness is a measure of the asymmetry of the distribution about its mean.
@@ 4.3.1 Negative Kurtosis Attack
The uniform distribution increased the relative certified radius the most and it builds the intuition behind the negative kurtosis attack. By modifying the distribution bits, so that they are no longer uniformly distributed, but positively skewed, the kurtosis in the normal distribution is reduced, as there are fewer random numbers near the mean, and more towards the tails. The modification to the PCG64 code is shown in LABEL:lst:pcg-kurtosis. Consider a uniformly distributed random variable, $X=\mathcal{U}(0,1)$. The probability density function (pdf) for $X$ is:
@@
This can be skewed to a random variable $X^{\prime}$ by altering the pdf as follows (for $b$ in Equation 14):
@@
$x^{\prime}$ $=\frac{-b+\sqrt{b^{2}-4\frac{ax}{2}}}{2\frac{a}{2}}$ (12)
$=\frac{-b+\sqrt{b^{2}-2ax}}{a}.$ (13)
@@
In LABEL:lst:pcg-kurtosis, the tunable parameter is $\alpha$, such that $a=1/\alpha$. First, the 64-bit random integer is converted to a 64-bit double precision floating point number. Then, the transformation from $X$ to $X^{\prime}$ is applied. Next, the number is converted back to a 64-bit integer. This is finally bit-shifted left by 9 bits and the least significant 9 bits are copied over from the original random integer generated by the PRNG. This is so that the index and sign bits remain random, and only the distribution bits are modified. A lower value of $\alpha$ results in higher skewness in the distribution bits, and more negative kurtosis in the resulting distribution. The sampled probability densities are plotted in Figure 1(a), along with a reference curve of the probability density function of the normal distribution.
@@ 4.3.2 Skewness Attack
The absolute normal distribution achieved a significant success rate in manipulating the certified radius to be high enough such that evasion attacks could be successful, and yet still maintain an acceptable level of performance compared to the original model. Therefore, this distribution formed the basis of the skewness attack as shown in LABEL:lst:pcg-skew. An 8-bit integer is used as a counter, initialised to 1. Every time a random integer with the sign bit set to 1 is encountered, the counter is incremented. When the value of the counter reaches the tunable parameter, $\beta$, the sign bit is flipped to zero, keeping the rest of the random number the same. This results in the final normally distributed random number being positive instead of negative, thereby skewing the distribution. The sampled probability distributions for $\beta=0,1,2,$ and $4$ are shown in Figure 1(b). For $\beta=0$, the resulting distribution is the same as $|\mathcal{N}(0,0.25^{2})|$.
@@ 4.3.3 Positive Kurtosis Attack
The motivation for this attack comes from the increase in certified radius observed when the normal distribution was replaced with the Laplace distribution. While the relative increase in radius achieved was not as high as the others, the attack is further motivated by the reasoning that it is more likely that the sampled prediction is the same as the modal class if the l2 norm of the additive noise is lower, thereby increasing the certified radius. The modified function in PCG for this attack is shown in LABEL:lst:pcg-zero.
A counter is initialised to 1 at first and is incremented by 1 every time a random number is generated. Every time the value of the counter is equal to the tunable parameter $\gamma$, the index bits in the random number are shifted right by 1. This divides the value of the index and reduces the size of the rectangular region chosen by the algorithm used for transformation to the normal distribution. It moves the random normal numbers closer to the mean, and increases the kurtosis of the resulting distribution, with the maximum kurtosis achieved for $\gamma=0$, and becoming closer to the standard normal distribution as $\gamma$ increases. The sampled probability densities of the distributions resulting from $\gamma=0,1,2$ and $4$ are shown in Figure 1(c).
@@ 4.4.1 NIST Test Suite
The National Institute of Standard and Technology (NIST) first published the SP800-22 test suite in 2001. Its purpose is to to ensure that random number generators used for cryptography are secure. This is done via a set of 15 statistical tests that operate on bitstreams from a random number generator, aiming to determine any deviation from a truly random sequence of bits. Brief descriptions of all 15 tests are given in Appendix C, including recommendations for overall size of the bitstream and test parameter choice. The reader is directed to Bassham et al. for a more in-depth description [5].
@@ 4.4.2 Test for Normality
We employ a number of statistical tests to detect finer fluctuations in the resulting distribution after transformation, and to quantify the deviation. In the main body of the paper we use the Shapiro-Wilk test – empirically one of the highest-power statistical tests to determine whether a set of samples is normally distributed [55]. We list performance of other tests in Appendix F. The test statistic for this test [61] is defined as:
@@ 5.1 Experimental Setting
Setting: The certification is performed on a subsample of 500 images from the CIFAR 10 test set [35], each with 100 noise samples for selection ($n_{0}$), 10,000 noise samples for certification ($n$), and $\alpha=0.001$. We use the same base model as Cohen et al. [13], with a ResNet-110 architecture.
@@ 5.2 Naive attacker
Table 2: This table shows the relative certified radius, $R^{\prime}/R$, where $R^{\prime}$ is the manipulated radius under attack, and $R$ is the baseline radius. The fraction of images for which $R^{\prime}/R$ falls in different bins is shown for the naive attacks and all three of the bit-flipping PRNG attacks. The maximum value of the relative certified radius achieved for each attack is also reported. The values in red indicate instances when the attack managed to successfully manipulate the radius by a factor of at least 1.5. Out of the naive attacks, $\mathcal{U}(-0.25,0.25)$ achieves the best performance, followed by $|\mathcal{N}(0,0.25^{2})|$, $\mathcal{L}(0,0.25)$, and finally $\mathcal{B}(0.5)$. Among the bit-flipping PRNG attacks, the skewness performs the best, then the positive-kurtosis, followed by the negative kurtosis.
@@
Table 3: This table shows the results of the NIST test suite for cryptographically secure RNGs on the three bit-flipping PRNG attacks with varying values of $\alpha$, $\beta$ and $\gamma$. The results for 4 popular modern PRNGs, MT19337, PCG64, Philox and SFC64 are also shown. Tests were run for bit streams of length $10^{6}$ and the reported values are the number of instances of each test that passed out of 1000. The minimum pass rate for a good RNG recommended by NIST is 980. The frequency, cumulative sums and runs tests managed to detect all attacks with high confidence.

### v2

@@ 3.4 Attacks, Dual_EC_DRBG and Bullrun
Traditional attacks on pseudo-random number generators have focused on determining the next number to be generated in a sequence; or more generally, given some of the numbers in a sequence, deducing some others. If the generator is used to produce cryptographic keys and also other random values that may become visible to the attacker, such a deduction may compromise an encrypted message. Such attacks aim to determine the state of the PRNG being used by the victim. Since most PRNGs are deterministic, a known state can be used to determine future numbers in the sequence [20]. Another class of attacks involves introducing a backdoor in the PRNG to make future outputs more predictable. The NSA pushed NIST to include EC_DRBG in the ANSI X9.82 standard for random number generators when the standard was first developed at the start of the 21st century. The PRNG featured in the first draft of the standard, which was published in 2004. The research community expressed concerns about EC_DRBG not being cryptographically sound by 2005 [23] – before the first official version of the standard was published [4] – leading to conjecture of an NSA backdoor [58]. Despite these concerns, EC-DBRG was adopted as the default random number generator by RSA in their BSAFE cryptography library [14]. The Snowden leaks in 2013 confirmed the existence of project Bullrun, which aimed “to covertly introduce weaknesses into the encryption standards followed by hardware and software developers worldwide" and successfully injected a backdoor into the default PRNG used for all cryptographic encryption between 2006 and 2013. This is a useful reminder of the tactics, tools, and procedures that may be used by a capable motivated adversary to carry out a backdoor attack against pseudo-random number generators.
@@ 4.1 Threat Model
It is assumed that the victim is attempting to use ML to get predictions in a safety-sensitive or security-sensitive setting. Therefore it is important to accurately gauge the robustness of every prediction, and we assume that Randomised Smoothing is being employed for this purpose. By definition, the probability that an adversarial example can be found within the certified radius obtained from randomised smoothing is low: less than 0.1% if $\alpha$ is set to $0.001$ as suggested by Cohen et al. [12]. The knowledge that the victim is employing randomised smoothing can in itself be very useful to an adversary. This was demonstrated by Cullen et al. [15], who used this information to only search for adversarial examples with $l^{2}$ distance greater than the certified radius, achieving better than the state-of-the-art success rate at finding adversarial examples against ML models using randomised smoothing. Furthermore, in their evaluation of randomised smoothing, Cohen et al. [12] found that the probability of finding adversarial examples increases rapidly as the upper bound on the $l^{2}$ norm of the adversarial example set by the adversary increases beyond the certified radius, $R$. For the ImageNet dataset [16], they found that the probability of finding an adversarial example against a smoothing classifier is 0.17 at an upper bound of $1.5R$ and 0.53 at $2R$. Hence, confidence serves a good proxy for reliability of prediction, and any compromise of Randomised Smoothing will decrease its trustworthiness.
@@
The objectives of an attacker attempting the class of attacks presented in this paper are twofold, depending on whether the certified radius obtained from randomised smoothing is being manipulated to be higher or lower. A higher certified radius can make the victim believe that the prediction is more robust for a given input than it actually is, so that the victim ignores adversarial examples within the spoofed radius. And certification is costly, requiring approximately $10^{5}$ inferences from the smoothed model to certify a radius of $4\sigma$. Reducing the certified radius can make the victim waste time and compute, forcing them to generate more predictions under noise to obtain the desired radius, providing a service-denial attack.
In this paper we limit our adversary access to the underlying random number generator used by the victim to perform randomised smoothing. While the setting may seem unrealistic, an attack with such access happened in the past. The story of the Dual_EC_DBRG PRNG, which the NSA used to compromise RSA libraries, was told in Section 3.4.
@@
The setup used by the victim to perform training and certification can be quite complicated, leaving multiple avenues for an attacker to gain access and manipulate the noise. Modern practices such as ML-as-a-service, and outsourcing of training and data generation to third-parties, opened a Pandora’s box. The attacks discussed here focus on modifying the noise distribution, first by using a different noise function directly and later by modifying the bitstream generated by an underlying pseudo-random number generator (PRNG). Where the victim has little or no control over the software and hardware used for training and certification, there are even more attack vectors. The objective of the attacks we present in the following sections is to spoof the certified radius while escaping detection by the victim – whether by analysing the performance of the model or by using statistical tests (discussed in Section 4.4). This class of attacks can be carried out in a traditional way by exploiting the hardware or software layers. In addition, with recent developments in denoising diffusion models [26], they can be carried out in an ML-as-a-Service scenario too. If the victim submits inputs to a cloud API for prediction and certification, an attacker can effectively remove the noise introduced by the victim and replace it.
@@ 4.2 Naive noise distribution replacement
In order to check the feasibility of the attacks described in the previous section, an initial test was done by explicitly changing the distribution that is sampled during certification. The GitHub repository released by Cohen et al. [^5] https://github.com/Xzh0u/randomized-smoothing was used as a starting point. Cohen et al. [12] reported that the variance of distribution that additive noise is sampled from controls the trade-off between accuracy and robustness. A lower value of $\sigma$ can be used to certify smaller radii but with a high degree of accuracy, whereas a higher value of $\sigma$ is needed to certify large radii, resulting in lower accuracy. $\sigma=0.25$ was found to achieve an acceptable trade-off between certification radius and accuracy on the CIFAR 10 [35] dataset, which was used to run all the experiments in this paper. Using a base classifier trained with additive noise sampled from $\mathcal{N}(0,0.25^{2})$, the naive attack swaps the noise distribution for one of the following when certification is performed: Laplace distribution, $\mathcal{L}(0,0.25)$; half-normal distribution, $|\mathcal{N}(0,0.25^{2})|$; uniform distribution, $\mathcal{U}(-0.25,0.25)$; and Bernoulli distribution, $\mathcal{B}(0.5)$. This naive attack was used to demonstrate the feasibility of this class of attacks on PRNGs. The next section will describe a more sophisticated attack.
Example Setting: Detection of an enemy tank is being performed from a satellite image. To make sure that the prediction is not a spurious correlation and to counteract camouflage paint, Randomised Smoothing is used. PRNG generates the noise for Randomised Smoothing. Randomness can come from the inference platform i.e. the cloud hardware or ML-as-a-Service API; alternatively the user can apply noise themselves [^6] Note that this may result in quantisation artefacts, e.g. pixel with value 245 noised to 235.3, gets quantised to 236, and potentially change performance of Randomised Smoothing. . A Normal User: A user attempting to certify the robust $l^{2}$ radius that does not contain adversarial examples around an input in the above setting. Here, our user may be an analyst that aims to find enemy tanks to target by a missile system. Attacker: An adversary with the objective of manipulating the certified radius obtained by the user. An attacker may increase it to make the prediction appear more robust than it actually is, i.e. convince an the user that tank is present at a given location, or decrease it to make the user uncertain about their prediction or even abstain from it, i.e. force them to not notice a tank at a given location. Defender: The defender’s goal is to detect the presence of a randomness-based adversary to determine if the noise function used for certification deviates significantly from white noise or not.
@@ 4.3 Bit-flipping PRNG attacker
The objective of the bit-flipping PRNG attacker is to modify the stream of bits being generated by the random number generator to alter the distribution when normal variates are sampled. An overview of the algorithm used to transform random integers to floats in the standard normal distribution is presented in Section 3.3. Taking a 64-bit random integer, the rightmost 8 bits determine the rectangle in which the normally distributed random number will fall. These are called the index bits. The 9th bit is the sign bit. Finally, the remaining 55 bits are transformed into a floating point number falling within the limits of the uniformly distributed rectangle chosen by the index bits. These are called the distribution bits. The following attacks modify one of these three categories of bits to alter the resulting distribution. Altering a distribution can be done by either introducing kurtosis or skewness. Kurtosis is a measure of the tailedness whereas skewness is a measure of the asymmetry of the distribution around the mean.
@@ 4.3.1 Negative Kurtosis Attack
The uniform distribution increased the relative certified radius the most and it builds the intuition behind the negative kurtosis attack. By modifying the distribution bits, so that they are no longer uniformly distributed, but positively skewed, the kurtosis in the normal distribution is reduced, as there are fewer random numbers near the mean and more towards the tails. The modification to the PCG64 code is shown in LABEL:lst:pcg-kurtosis. Consider a uniformly distributed random variable, $X=\mathcal{U}(0,1)$. The probability density function (pdf) for $X$ is:
@@
This can be skewed to a new random variable $X^{\prime}$ by altering the pdf as follows (for $b$ in Equation 11):
@@
$x^{\prime}=F^{\prime}(x)=\frac{-b+\sqrt{b^{2}-2ax}}{a}.$ (10)
@@
In LABEL:lst:pcg-kurtosis, the tunable parameter is $\alpha$, such that $a=1/\alpha$. First, the 64-bit random integer is converted to a 64-bit double-precision floating point number. Then, the transformation from $X$ to $X^{\prime}$ is applied. Next, the number is converted back to a 64-bit integer. This is finally bit-shifted left by 9 bits and the least significant 9 bits are copied over from the original random integer generated by the PRNG. This is so that the index and sign bits remain random, and only the distribution bits are modified. A lower value of $\alpha$ results in higher skewness in the distribution bits, and more negative kurtosis in the resulting distribution. The sampled probability densities are plotted in Figure 2(a), along with a reference curve of the probability density function of the normal distribution.
@@ 4.3.2 Skewness Attack
The half-normal distribution achieved a significant success rate in manipulating the certified radius to be high enough such that evasion attacks could succeed, and yet still maintain an acceptable level of performance compared to the original model. Therefore, this distribution formed the basis of the skewness attack shown in LABEL:lst:pcg-skew. An 8-bit integer is used as a counter, initialised to 1. Every time a random integer with the sign bit set to 1 is encountered, the counter is shifted left. When the value of the counter reaches the tunable parameter, $\beta$, the sign bit is flipped to zero, keeping the rest of the random number the same. This results in the final normally distributed random number being positive instead of negative, thereby skewing the distribution. The sampled probability distributions for $\beta=0,1,2,$ and $4$ are shown in Figure 2(b). For $\beta=0$, the resulting distribution is the same as $|\mathcal{N}(0,0.25^{2})|$.
@@ 4.3.3 Positive Kurtosis Attack
The motivation for this attack comes from the increase in certified radius observed when the normal distribution was replaced with the Laplace distribution. While the relative increase in radius achieved was not as high as the others, the attack is further motivated by the reasoning that it is more likely that the sampled prediction is the same as the modal class if the $l^{2}$ norm of the additive noise is lower, thereby increasing the certified radius. The modified function in PCG for this attack is shown in LABEL:lst:pcg-zero.
A counter is initialised to 1 and shifted left every time a random number is generated. Every time the value of the counter is equal to the tunable parameter $\gamma$, the index bits in the random number are shifted right by 1. This divides the value of the index and reduces the size of the rectangular region chosen by the algorithm used for transformation to the normal distribution. It moves the random normal numbers closer to the mean, and increases the kurtosis of the resulting distribution, with the maximum kurtosis achieved for $\gamma=0$, and becoming closer to the standard normal distribution as $\gamma$ increases. The sampled probability densities of the distributions resulting from $\gamma=0,1,2$ and $4$ are shown in Figure 2(c).
@@ 4.4.1 NIST Test Suite
The National Institute of Standards and Technology (NIST) first published the SP800-22 test suite in 2001, in the aftermath of the competition in which the AES algorithm was selected. Its purpose was to ensure that random number generators used for cryptography are secure. This is done via a set of 15 statistical tests that operate on bitstreams from a random number generator, aiming to detect any deviation from a truly random sequence of bits. The reader is directed to Bassham et al. for an in-depth description of each test along with the recommended parameter values to use [5]. It is important to note that this test suite is a measure of PRNG badness and not goodness. According to the recommendations presented by NIST, it is important that a PRNG’s design is based on sound mathematical principles. Due to the statistical nature of these tests, it is possible that a well-designed PRNG may sometimes fail a test, while a bad PRNG may pass some tests, as demonstrated later in this paper.
@@ 4.4.2 Test for Normality
We employ a number of statistical tests to detect finer fluctuations in the resulting distribution after transformation and to quantify the deviation. In the main body of the paper, we use the Shapiro-Wilk test [59] – empirically one of the highest-power statistical tests to determine whether a set of samples is normally distributed [53]. We list the performance of other tests in Appendix D. The test statistic for this test is defined as:
@@ 5.1 Experimental Setting
Setting: Certification is performed on a subsample of 500 images from the CIFAR 10 test set [35], each with 100 noise samples for selection ($n_{0}$), 10,000 noise samples for certification ($n$), and $\alpha=0.001$. We use the same base model as Cohen et al. [12], with a ResNet-110 architecture.
@@ 5.2 Naive attacker
Table 2: This table shows the relative certified radius, $R^{\prime}/R$, where $R^{\prime}$ is the manipulated radius under attack, and $R$ is the baseline radius. The fraction of images for which $R^{\prime}/R$ falls in different bins is shown for the naive attacks and all three of the bit-flipping PRNG attacks. The maximum value of the relative certified radius achieved for each attack is also reported. The values in red indicate instances when the attack managed to successfully manipulate the radius by a factor of at least 1.5. Out of the naive attacks, $\mathcal{U}(-0.25,0.25)$ achieves the best performance, followed by $|\mathcal{N}(0,0.25^{2})|$, $\mathcal{L}(0,0.25)$, and finally $\mathcal{B}(0.5)$. Among the bit-flipping PRNG attacks, the skewness performs the best, then the positive-kurtosis, followed by the negative-kurtosis.
@@
Table 3: This table shows the results of the NIST test suite for cryptographically secure PRNGs on the three bit-flipping PRNG attacks with varying values of $\alpha$, $\beta$ and $\gamma$. The results for 4 popular modern PRNGs, MT19337, PCG64, Philox and SFC64 are also shown. PCG64 and SFC64 are NIST certified cryptographically secure PRNGs, MT19337 is the default PRNG in Python and Philox is the default PRNG in PyTorch. Tests were run for bit streams of length $10^{6}$ and the reported values are the number of instances of each test that passed out of 1000. The minimum pass rate for a good RNG recommended by NIST is 980. The frequency, cumulative sums and runs tests managed to detect all attacks with high confidence.